OpenStudy (alexh107):

Simplify ((sqrt5) -2i)^2

OpenStudy (alexh107):

It looks like this: \[(\sqrt{5}-2i)^2\]

OpenStudy (sshayer):

\[\left( a-b \right)^2=a^2+b^2-2ab\]

OpenStudy (learner):

In general, \(\color{black}{\displaystyle (a+bi)^2=a^2+2a(bi)+(bi)^2=a^2+2abi+(b^2)(i^2)=a^2+2abi-b^2}\).

OpenStudy (alexh107):

I had it set up as (\[(\sqrt{5}-2i)(\sqrt{5}-2i)\] and then I started multiplying it out and got stuck after \[\sqrt{25}\]

OpenStudy (learner):

Give your best shot, and we will fix your work if anything ...

OpenStudy (alexh107):

I'm not sure the rules for multiplying \[\sqrt{5} and -2i\]

OpenStudy (sshayer):

\[\sqrt{5}*\sqrt{5}=\left( \sqrt{5} \right)^2=5\] or \[\sqrt{25}=5\]

OpenStudy (sshayer):

you are doing good.

OpenStudy (learner):

\(\color{black}{\displaystyle (\color{red}{\displaystyle \sqrt{{\tiny~}5{\tiny~~}}}-\color{blue}{2i})^2=(\color{red}{\sqrt{{\tiny~}5{\tiny~~}}})^2-2(\color{red}{\sqrt{{\tiny~}5{\tiny~~}}})(\color{blue}{2i})+(\color{blue}{2i})^2}\)

OpenStudy (learner):

sshayer mentioned this rule in his first reply

OpenStudy (learner):

the first two terms shouldn't be hard, and as regards to the last term, the last term is in a form \((bi)^2\), and this is how it is simplified: \((bi)^2=b^2\times i^2=b^2\times (-1)\) (since \(i^2=-1\)) \(=-b^2\).

OpenStudy (phi):

*** I'm not sure the rules for multiplying \(\sqrt{5}\) and−2i *** first, remember that -2i means -2 * i so you multiply all three to get \[ \sqrt{5} \cdot -2 \cdot i \] usually, people write the -2 first, and leave out the multiply signs \[ -2 \sqrt{5} i \] but it means the same thing : -2* sqr(5) * i